# Certified spectral approximation of transfer operators and the Gauss map

### v1

@@ Abstract
We prove that the full spectral picture of a transfer operator (every isolated eigenvalue, eigenvector, and Riesz projector outside the essential spectral radius) can be approximated to arbitrary precision by finite-rank discretizations, with no spectral pollution. The method is a posteriori: once a computable approximation bound is available (from compactness or a Doeblin–Fortet–Lasota–Yorke inequality, which may require hyperbolicity constants and adapted Banach spaces), the spectral gap and multiplicity are certified from computed data via a resolvent perturbation bound, with no further dynamical input. This applies both to compact operators on a single Banach space and to quasi-compact operators satisfying a Doeblin–Fortet–Lasota–Yorke inequality, extending Li’s resolution of the Ulam conjecture from the invariant density to the entire discrete spectrum with certified error bounds at every finite truncation.
As a benchmark, we certify the first $50$ nonzero eigenvalues of the Gauss–Kuzmin–Wirsing operator to at least $90$ rigorous decimal digits, together with their eigenvectors, Riesz projectors, and spectral gap, yielding a certified spectral expansion for Gauss–Kuzmin distributions with explicit error bounds and providing a rigorous answer to the Gauss–Babenko–Knuth problem on the spectral data of the Gauss map.
@@ 1 Introduction and overview
Transfer operators encode invariant measures, rates of mixing, limit theorems, and Ruelle–Pollicott resonances, yet extracting quantitative spectral information from an infinite-dimensional operator remains delicate: one must control discretization/truncation effects, resolvent growth, and spectral-projector stability with fully explicit constants.
The central message of this paper is that, under broadly available regularization hypotheses (compactness/nuclearity on a suitable space, or a Lasota–Yorke inequality), the spectral analysis of an infinite-dimensional operator can be reduced to a finite-dimensional, computationally certifiable task. Starting from a finite-rank discretization $L_{N}$ of an operator $L$ with an explicit error bound, we compute spectral objects for $L_{N}$ (eigenvalues, resolvent bounds, Riesz projectors) and transfer them to $L$ via a resolvent perturbation bound with fully explicit constants. The analytic work is confined to producing a few quantitative constants, after which the spectral certification becomes validated numerical linear algebra. The a posteriori approach to certifying spectral data of transfer operators from computed information originates in [28, 27, 29, 26, 25], which developed rigorous certification of invariant measures, mixing rates, and escape rates via Lasota–Yorke inequalities and the coarse-fine approximation strategy. [12] extended the approach to eigenvalue enclosures and Riesz projector certification via validated linear algebra. Here we isolate a set of hypotheses applicable in a range of functional-analytic settings, extend the framework to the Doeblin–Fortet–Lasota–Yorke (DFLY) setting, and develop high-precision implementations that certify many eigenvalues, eigenvectors, and Riesz projectors simultaneously.
@@
Computability of isolated spectral data.
A longstanding open problem, raised by Arveson [2] and Böttcher [13], asks whether the spectrum of a bounded operator can be computed from its finite-dimensional sections. The Solvability Complexity Index (SCI) hierarchy developed by Colbrook, Hansen, and collaborators [10, 17, 18] resolves this question by showing that, for general bounded operators on a separable Hilbert space, computing the spectrum requires several successive limits and no single algorithm can provide error control at any finite step [10, Theorem 1.1]. Even for compact operators the obstruction persists: convergent algorithms exist, but no algorithm that takes only matrix entries as input can certify that its output is within a prescribed distance of the true spectrum.
@@ Acknowledgements
The author would like to acknowledge helpful discussions with A. Blumenthal, O. Bandtlow, and S. Galatolo. The idea of treating the Gauss map arose during a discussion with I. Melbourne. The author is also grateful to the participants of related seminars and workshops for valuable feedback on preliminary versions of this work.

### v2

# arXiv:2602.19435v2[math.DS] 01 Mar 2026

@@ Abstract
Existing rigorous frameworks for transfer operators require spectral gaps as a priori input. We reverse the logical direction: spectral gaps, eigenvalue multiplicities, and projector stability are certified a posteriori from finite-rank computed data, with no spectral assumption beyond a computable approximation bound.
The method applies to compact operators on a single Banach space and to quasi-compact operators satisfying a Doeblin–Fortet–Lasota–Yorke inequality, extending Li’s resolution of the Ulam conjecture from the invariant density to the entire discrete spectrum with certified error bounds at every finite truncation. For the discrete spectrum of transfer operators, this resolves a computability obstruction identified in the Solvability Complexity Index hierarchy: the dynamical structure furnishing the truncation bound is precisely the oracle that breaks the barrier preventing spectral certification from finite-dimensional sections.
As a benchmark, we certify the first $50$ nonzero eigenvalues of the Gauss–Kuzmin–Wirsing operator to at least $90$ rigorous decimal digits, together with eigenvectors, Riesz projectors, and spectral gap, providing a rigorous answer to the Gauss–Babenko–Knuth problem.
@@ 1 Introduction and overview
Transfer operators govern invariant measures, rates of mixing, and limit theorems through their spectral data (eigenvalues, called Ruelle–Pollicott resonances in this context, and eigenprojectors), yet extracting quantitative spectral information from an infinite-dimensional operator remains delicate: one must control discretization/truncation effects, resolvent growth, and spectral-projector stability with fully explicit constants.
The central message of this paper is that, under broadly available regularization hypotheses (compactness/nuclearity on a suitable space, or a Lasota–Yorke inequality), the spectral analysis of an infinite-dimensional operator can be reduced to a finite-dimensional, computationally certifiable task. Starting from a finite-rank discretization $L_{N}$ of an operator $L$ with an explicit error bound, we compute spectral objects for $L_{N}$ (eigenvalues, resolvent bounds, Riesz projectors) and transfer them to $L$ via a resolvent perturbation bound with fully explicit constants. The analytic work is confined to producing a few quantitative constants, after which the spectral certification becomes validated numerical linear algebra. The a posteriori approach to certifying spectral data of transfer operators from computed information originates in [28, 27, 29, 26, 25], which developed rigorous certification of invariant measures, mixing rates, and escape rates via Lasota–Yorke inequalities and the coarse-fine approximation strategy. A subsequent work [12] extended the approach to eigenvalue enclosures via validated linear algebra. Here we isolate a set of hypotheses applicable in a range of functional-analytic settings, extend the framework to the Doeblin–Fortet–Lasota–Yorke (DFLY) setting, and develop high-precision implementations that certify many eigenvalues, eigenvectors, and Riesz projectors simultaneously.
@@
Computability of isolated spectral data.
A longstanding open problem, raised by Arveson [2] and Böttcher [13], asks whether the spectrum of a bounded operator can be computed from its finite-dimensional sections. The Solvability Complexity Index (SCI) hierarchy developed by Colbrook, Hansen, and collaborators [10, 17, 18] resolves this question by showing that, for general bounded operators on a separable Hilbert space, computing the spectrum requires several successive limits and no single algorithm can provide error control at any finite step. In particular, computing the discrete spectrum and determining spectral gaps are both $\Sigma_{2}^{A}$-complete: they require two successive limits, and no algorithm taking only matrix entries as input can certify that its output is within a prescribed distance of the true spectral data [17, Theorems 3.13 and 3.15].
@@ Acknowledgements
The author would like to acknowledge helpful discussions with A. Blumenthal, O. Bandtlow, S. Galatolo, and W. Tucker. The idea of treating the Gauss map arose during a discussion with I. Melbourne. The author is also grateful to the participants of related seminars and workshops for valuable feedback on preliminary versions of this work.